A fluidization catalyst, specific gravity 1.8, is fluidized with air at 550K and 2atm. The catalyst may be assumed spherical, with diameter 300 um. The gravitational acceleration is 9.8 ms-2. The static void fraction is 0.42. The unfluidized bed height is 1.5 m. Find in SI units: The minimum pressure drop will be: *
Answer #1

g lm3 -1.2838 74 Pa = 1283-874 Pa kg/m3 [as 1 kg = 1 oso g Now, minimen pressure doop (OP min) is given by Almin =g (P.-Pa) (